# X-ray Variability in the Symbiotic Binary RT Cru: Principal Component Analysis

### v1

@@ 4.1 Computational Approach
Principal Components Analysis (PCA), which is a highly effective technique for deconstructing temporal variable data, has frequently been used in the field of multivariate statistical research in astronomy, such as gaining insights into the spectral types of stars (Deeming, 1964; Whitney, 1983), statistical analysis of galaxies (Bujarrabal et al., 1981; Efstathiou & Fall, 1984), multiple epoch UV data of active galaxies (Mittaz et al., 1990), distinct emission components in optical observations of quasars (Francis et al., 1992; Boroson & Green, 1992), and imaging analysis of molecular clouds (Heyer & Schloerb, 1997; Brunt & Heyer, 2002). This eigenvector-based multivariate method was later employed to analyze the X-ray spectral variability of Seyfert galaxies (Vaughan & Fabian, 2004; Miller et al., 2007; Parker et al., 2014a; Gallo et al., 2015) and X-ray binaries (Malzac et al., 2006; Koljonen et al., 2013; Koljonen, 2015). The PCA mathematical framework for spectral analysis is similar to that adopted in the flux-flux correlation method as shown by Vaughan & Fabian (2004), which results in multiple spectrally-invariant components whose amplitudes exhibit temporal variation. The observed variable spectra can be regenerated by linearly combining all these PCA components. In theory, this approach should enable us to identify all the variable components of any given multidimensional dataset. However, the assumption of spectrally-invariant components may not hold true in real observational data, where noise in the background is likely to cause some variations. As demonstrated by Parker et al. (2015), one of the key benefits of employing multivariate analysis is the ability to determine the minimum number of PCA components required for building the observed variations, which may not be created by noise fluctuations. A clear result of PCA would be time-dependent emission or absorption components with little variability but enough statistical significance to not be distinguished as background noise (see, e.g., Koljonen et al., 2013; Parker et al., 2015, 2017).
@@
To conduct our principal component analysis of RT Cru, we used a customized implementation of the Python program pca originally written by Parker et al. (2018). This program leveraged the singular value decomposition (svd) function (Press et al., 1997) from the linear algebra (linalg) submodule of the Python library NumPy. It also utilizes the effective area column read from the ARF data to convert the count spectrum to the photon-flux one , i.e., , where is the effective function, and the exposure time. The time-sliced spectrum of the background is also used to eliminate background contamination from the source. Moreover, a mean spectrum, calculated with the time-sliced spectra of each dataset was used to derive the normalized time-sliced spectra as follows . The purpose of this program is to transfer a set of time-sliced spectra binned at energy intervals into a 2D array (). The resulting array is then subjected to a decomposition process using the svd function, resulting in a matrix () including the principal components sorted with their eigenvalues, in addition to an -array of eigenvalues yielding the fractional variability, as well as a matrix () containing the eigenvectors representing the time series . The PCA spectra and their corresponding light curves describe the spectral characteristics and their temporal variations present in a complex variable source, respectively. The fractional variability of each component is estimated using the normalized eigenvalues. The uncertainty calculation in the PCA program is implemented according to the procedure outlined in Miller et al. (2007), where the spectra are subjected to random perturbations and subsequent recalculations with the svd function.
@@ 4.2 PCA Results
Figure 3 shows the spectrum and its time series of the first components produced from the Chandra HRC-S/LETG (top panels) and XMM-Newton EPIC-pn data (middle panels), along with the corresponding log-eigenvalue (LEV) diagrams (right). We should note that the number of non-zero eigenvalues provided by the svd function corresponds to the number () of time-segmented PHA spectra in each LEV diagram. The black line in each LEV diagram corresponds to a linear correlation found between logarithmic normalized eigenvalues and eigenvector orders of the components with orders higher than 3, which should be associated with noise rather than real spectral variations. The LEV diagrams help us determine those components that are statistically significant as described by Parker et al. (2018). It can be seen that the first component in each dataset, whose spectrum and time series are plotted, are slightly above the high-order correlation line. The peak and valley features that appear in the PCA spectra could be indicators of emission and/or absorption lines (see, e.g., Parker et al., 2017, 2018). We see some spectral features in the soft band, which might suggest the presence of emission lines from H-like and He-like ions, N vii Ly 0.5 keV, N vi He 0.43 keV, O viii Ly 0.65 keV, and O vii He 0.57 keV. However, it is difficult to detect these thermal emission features in the soft excess, even if they do exist, because they are mixed up with high background noise. A recent investigation by Zhang et al. (2023) could not constrain these emission lines but put an upper confidence level of 1 keV on the temperature of a soft thermal plasma component. The time series of the first components suggest that these emission lines, even if they are real, appear temporarily during brightening events occurring every 20-25 ks and lasting for ks, which makes it difficult to constrain them. Although the effectiveness of the Chandra HRC-S/LETG in the hard band is low, the XMM-Newton EPIC-pn instrument seems to capture some weak features in the energies associated with Ar xviii Ly 3.3 keV, Ca K 4 keV, and Fe K 6.4 keV. Nevertheless, the emission lines look to be more predominant in the soft band ( keV) compared to the hard band in the spectrum of the XMM-Newton data.
@@
In Fig. 3, we also present the spectra and the corresponding light curves of the first three PCA components generated from the Suzaku XIS1 data. These three components are statistically significant according to their noticeable deviations from the high-order eigenvalue regression in the LEV diagram. As seen in Fig. 2, the source exhibited increases in both the softness (lower ) and the hardness (higher ) in 2012 compared to 2007, while the brightness in the total band () was lower in 2012 compared to that in 2007.
@@
The first component in the Suzaku XIS1 data is akin to the spectral features of an absorbing column component. The second component shows a blackbody-like emission below 2 keV followed by an absorbed hard continuum above 2 keV, but with a peak at keV, corresponding to the iron K line. The first two components depicts a highly absorbed continuum source, which is similar to the model used by Danehkar et al. (2021). Additionally, our multivariate analysis suggests a (multi-) blackbody-like thermal component in the soft excess in the Suzaku observations. The normalized eigenvalues yield the variability fractions of % and 20% for the spectra and , respectively. This implies that the variability in the absorbing material and the source continuum are mainly responsible for the changes in the hardness ratios and between 2007 and 2012. This phenomenon might be related to the almost disappearance of the hard X-rays identified in 2019, which was argued to be associated with a substantial reduction in the falling material (Pujol et al., 2023). In addition, the absorbing column contributes about twice the dramatic changes in the hardness ratios of the source continuum over the two epochs. The light curves and depict that as the source got fainter and harder in the last epoch compared to the prior epoch, it had higher obstruction in 2012 compared to 2007.
@@
The spectrum and its light curve of the first PCA component extracted from the Suzaku HXD-PIN data are shown in Fig. 4, along with the log-eigenvalue diagram. The comparison between the time series of the XIS1 and PIN data indicates that both of them have similar changes over time, so the PCA spectrum derived from the PIN observations corresponds to the extreme hard excess of the continuum seen in of the XIS1 data. However, there is no indication of the absorbing column in the eigenvector-based multivariate analysis of the PIN data. This could mean that the absorbing material mostly blocks the energy range below 5 keV, as seen in of the XIS1 in Fig. 3. In addition, there is no PCA component containing the thermal emission lines similar to the spectrum of the XIS1, so they could be the features predominately present in the soft excess. We should also note that the PIN has an energy resolution of around 3 keV in the 10-30 keV range, which is lower than the Suzaku XIS (120 eV at 6 keV), so the PIN cannot resolve any emission line features.
Figure 4 also presents the spectra and the associated time series of the first two PCA components deduced from the NuSTAR observation. The first-order PCA spectrum shows a powerlaw-like continuum similar to that seen in of the Suzaku HXD-PIN observations. Additionally, the second PCA component may contain some emission features at energies typically associated with Ar xviii Ly 3.3 keV, Ca K 3.7 keV, and Ca K 4 keV, which are possibly present in of the XMM-Newton EPIC-pn data and of the Suzaku XIS1 observations. The light curve of the NuSTAR PCA suggests that the second component, which likely contains some line features, appears during brightening events occurring on an intervals between 30-40 ks. However, we caution that NuSTAR’s energy resolution of eV is lower than those of the XMM-Newon EPIC-pn (80 eV) and Suzaku XIS (50 eV at 1 keV). In addition, there is no evidence for the obscuring material in the multivariate statistical analysis of the NuSTAR data, implying the absorbing material is not largely variable during the course of the NuSTAR observation similar to those seen in the Suzaku observations.
@@ 4.3 Constructing Spectra from PCA
To better evaluate the nature of the PCA components from Suzaku/XIS1, we used their corresponding photon-flux data to create the xspec-compatible spectra and the response files using the ftool program ftflx2xsp. The photon-flux spectra were reconstructed according to , where is the normalized spectrum of the -order principal component produced by the svd function, is the mean spectrum derived from the time-sliced spectra, and is the normalized eigenvalue of the -order principal component created via svd, representing the variability fractions. We analyzed the reconstructed spectra in the Interactive Spectral Interpretation System (isis v 1.6.2-51; Houck & Denicola, 2000) that has access to the xspec models (Arnaud, 1996).
@@
Figure 5 shows the reconstructed spectra of the first and second principal components of the Suzaku XIS1 data, which were modeled using a phenomenological model, pcfabs (diskbb apec) pcfabs (compTT zgauss), consisting of an accretion disk model (diskbb), a collisionally-ionized diffuse model of the Astrophysical Plasma Emission Code (apec; Smith et al., 2001), a Comptonization model (compTT), partial covering fraction absorption components (pcfabs), and Gaussian components (zgauss). The soft excess was well reproduced using a diskbb model made of multiple blackbody components (see, e.g., Mitsuda et al., 1984; Makishima et al., 1986) and an emission spectrum produced by the collisional plasma apec model; both of them are partially covered by an absorbing column. To create phenomenologically the curvature in the hard excess, we employed a partially covered, absorbed Comptonization model (compTT) of the soft radiation in a hot plasma cloud analytically obtained by Titarchuk (1994) and Titarchuk & Lyubarskij (1995). To improve the model fit, three Gaussian components were also included, namely Fe K (6.379 keV), Fe He (6.693 keV), and Fe Ly (6.946 keV). We caution that the PCA components are built from the Suzaku data according to the temporal evolution over the two epochs, so not all the derived spectral features are physically real.
@@
Table 3 lists the best-fitting values of the parameters, obtained using the Levenberg-Marquardt minimization algorithm (Moré, 1978) and the chi-square statistic (; see Bevington & Robinson, 2003), with the uncertainties at 90% confidence derived using the isis standard function for confidence limits (conf_loop). It can be seen that the soft excess is partially covered by an absorbing column of cm-2. However, the covering fraction in the first PCA spectrum () is higher than that in the second PCA spectrum (). Moreover, the hard excess is partially absorbed by columns of and cm-2 in the first and second PCA components, respectively; both of them with a covering fraction of . This implies that there is more absorbing material in the first principal component. Intrestingly, Luna et al. (2018) also found that the absorbing column in 2012 is about cm-2 higher than that in 2007 based on the XIS1 data.
@@
In addition, the curvature in the soft excess was phenomenologically reproduced using a multiple-blackbody accretion disk with a temperature of keV at the inner radius, along with a collisionally-ionized soft thermal plasma emission with temperatures of 1.1 and 1.2 keV in the first and second PCA components, respectively. We should note that the inclusion of the apec collisional plasma model helps to recreate the PCA spectra, while it could not be done using only a diskbb model. In particular, Danehkar et al. (2021) previously deduced a heavily obscured, soft thermal plasma emission with a temperature of 1.3 keV using low-count Bayesian statistics. Moreover, the third PCA component of the Suzaku XIS1 data, along with the principal component of the XMM-Newton EPIC-pn data, suggest the possible presence of some emission features mostly at the soft excess, which could be associated with thermal features of a collisional ionized plasma. So, the soft thermal apec component of - keV that appears in the spectral analysis of the PCA components might be physically related to that phenomenon.
@@
The soft component with a temperature of keV, if it exists, could be produced by the expanding winds or jets, similar to what was proposed for CH Cyg (Ezuka et al., 1998). The predicted X-ray temperature of the shock-ionized plasma created by the collision of the winds with the ISM is , where is the wind velocity, is the mean molecular weight, and the hydrogen mass (e.g., Güdel & Nazé, 2009). To produce an X-ray thermal spectrum of 1.2 keV, a wind velocity of 1,000 km s-1 is necessary. This is in the range of those estimated for CH Cyg (Karovska et al., 2007, 2010), which is another -type symbiotic star containing soft thermal spectra of 0.2 and 0.7 keV (Ezuka et al., 1998).
@@ 5.1 Transient Nature
The hard X-ray source IGR J12349-6434 was first detected with the IBIS instrument aboard INTEGRAL in 2003–2004 (Chernyakova et al., 2005). Its association with RT Cru was suggested by Masetti et al. (2005) and confirmed by Swift observations (Tueller et al., 2005b). Although the source was at a level of mCrab in the 20–60 keV energy band during the first detection (Chernyakova et al., 2005), it was seen 3.4 times brighter at 13 mCrab in the 18–40 keV energy band in 2012 (Sguera et al., 2012), followed by a flux level of mCrab in 2015 (Sguera et al., 2015). The ASAS and AAVSO optical light curves also indicated that RT Cru became brighter from 13.5 to 11.3 mag between 1998 and 2001, then a gradual decrease in the brightness to 12.1 mag in 2006, and later an increase to 11.8 mag in 2009 and 11.5 mag in 2012, followed by a decrease to 12.6 mag in 2017 (Luna et al., 2018). The Swift/BAT survey revealed that the source reached its highest X-ray brightness between 2011 and 2012 coincident with the optical peak in 2012 (Luna et al., 2018). As seen in Fig. 1, the Suzaku hardness ratios exhibited harder X-rays in 2012 compared to 2007. According to the spectral fitting results by Luna et al. (2018), this transition could be related to a later increase in the column density of the absorbing material, as well as a rise in the X-ray brightness owing to higher accretion rates.

### v2

@@ 4.1 Computational Approach
Principal Components Analysis (PCA), which is a highly effective technique for deconstructing temporally variable data, has frequently been used in the field of multivariate statistical research in astronomy, such as gaining insights into the spectral types of stars (Deeming, 1964; Whitney, 1983), statistical analysis of galaxies (Bujarrabal et al., 1981; Efstathiou & Fall, 1984), multiple-epoch UV data of active galaxies (Mittaz et al., 1990), distinct emission components in optical observations of quasars (Francis et al., 1992; Boroson & Green, 1992), and imaging analysis of molecular clouds (Heyer & Schloerb, 1997; Brunt & Heyer, 2002). This eigenvector-based multivariate method was later employed to analyze the X-ray spectral variability of Seyfert galaxies (Vaughan & Fabian, 2004; Miller et al., 2007; Parker et al., 2014a; Gallo et al., 2015) and X-ray binaries (Malzac et al., 2006; Koljonen et al., 2013; Koljonen, 2015). The PCA mathematical framework for spectral analysis is similar to that adopted in the flux–flux correlation method as shown by Vaughan & Fabian (2004), which results in multiple spectrally invariant components whose amplitudes exhibit temporal variation. The observed variable spectra can be regenerated by linearly combining all these PCA components. In theory, this approach should enable us to identify all the variable components of any given multidimensional dataset. However, the assumption of spectrally invariant components may not hold true in real observational data, where noise in the background is likely to cause some variations. As demonstrated by Parker et al. (2015), one of the key benefits of employing multivariate analysis is the ability to determine the minimum number of PCA components required for building the observed variations, which may not be created by noise fluctuations. A clear result of PCA would be time-dependent emission or absorption components with little variability but enough statistical significance to not be distinguished as background noise (see, e.g., Koljonen et al., 2013; Parker et al., 2015, 2017).
@@
To conduct our principal component analysis of RT Cru, we used a customized implementation of the Python program pca originally written by Parker et al. (2018). This program leveraged the singular value decomposition (svd) function (Press et al., 1997) from the linear algebra (linalg) submodule of the Python library NumPy. It also utilizes the effective area column read from the ARF data to convert the count spectrum to the photon-flux one , i.e., , where is the effective function and the exposure time. The time-sliced spectrum of the background is also used to eliminate background contamination from the source. Moreover, a mean spectrum, calculated with the time-sliced spectra of each dataset was used to derive the normalized time-sliced spectra as follows: . The purpose of this program is to transfer a set of time-sliced spectra binned at energy intervals into a 2D array (). The resulting array is then subjected to a decomposition process using the svd function, resulting in a matrix () including the principal components sorted with their eigenvalues, in addition to an -array of eigenvalues yielding the fractional variability, as well as a matrix () containing the eigenvectors representing the time series . The PCA spectra and their corresponding light curves describe the spectral characteristics and their temporal variations present in a complex variable source, respectively. The fractional variability of each component is estimated using the normalized eigenvalues. The uncertainty calculation in the PCA program is implemented according to the procedure outlined in Miller et al. (2007), where the spectra are subjected to random perturbations and subsequent recalculations with the svd function.
@@ 4.2 PCA Results
Figure 3 shows the spectrum and its time series of the first components produced from the Chandra HRC-S/LETG (top panels) and XMM-Newton EPIC-pn data (middle panels), along with the corresponding log-eigenvalue (LEV) diagrams (right). We should note that the number of non-zero eigenvalues provided by the svd function corresponds to the number () of time-segmented pulse-height amplitude (PHA) spectra in each LEV diagram. The black line in each LEV diagram corresponds to a linear correlation found between logarithmic normalized eigenvalues and eigenvector orders of the components with orders higher than 3, which should be associated with noise rather than real spectral variations. The LEV diagrams help us determine those components that are statistically significant as described by Parker et al. (2018). It can be seen that the first component in each dataset, whose spectrum and time series are plotted, is slightly above the high-order correlation line. The peak and valley features that appear in the PCA spectra could be indicators of emission and/or absorption lines (see, e.g., Parker et al., 2017, 2018). We see some spectral features in the soft band, which might suggest the presence of emission lines from H-like and He-like ions, N vii Ly 0.5 keV, N vi He 0.43 keV, O viii Ly 0.65 keV, and O vii He 0.57 keV. However, it is difficult to detect these thermal emission features in the soft excess, even if they do exist, because they are mixed up with high background noise. A recent investigation by Zhang et al. (2023) could not constrain these emission lines but put an upper confidence level of 1 keV on the temperature of a soft thermal plasma component. The time series of the first components suggest that these emission lines, even if they are real, appear temporarily during brightening events that occur every 20-25 ks and last for ks, which makes it difficult to constrain them. Although the effectiveness of the Chandra HRC-S/LETG in the hard band is low, the XMM-Newton EPIC-pn instrument seems to capture some weak features in the energies associated with Ar xviii Ly 3.3 keV, Ca K 4 keV, and Fe K 6.4 keV. Nevertheless, the emission lines look to be more predominant in the soft band ( keV) than the hard band in the spectrum of the XMM-Newton data.
@@
In Fig. 3, we also present the spectra and the corresponding light curves of the first three PCA components generated from the Suzaku XIS1 data. These three components are statistically significant according to their noticeable deviations from the high-order eigenvalue regression in the LEV diagram. As seen in Fig. 2, the source exhibited increases in both the softness (lower ) and the hardness (higher ) in 2012 compared to 2007, while the brightness in the total band () was lower in 2012 than in 2007.
@@
The first component in the Suzaku XIS1 data is akin to the spectral features of an absorbing column component. The second component shows a blackbody-like emission below 2 keV followed by an absorbed hard continuum above 2 keV, but with a peak at keV, corresponding to the iron K line. The first two components depict a highly absorbed continuum source, which is similar to the model used by Danehkar et al. (2021). Additionally, our multivariate analysis suggests a (multi-) blackbody-like thermal component in the soft excess in the Suzaku observations. The normalized eigenvalues yield the variability fractions of % and 20% for the spectra and , respectively. This implies that the variabilities in the absorbing material and the source continuum are mainly responsible for the changes in the hardness ratios and between 2007 and 2012. This phenomenon might be related to the almost total disappearance of the hard X-rays identified in 2019, which was argued to be associated with a substantial reduction in the falling material (Pujol et al., 2023). In addition, the absorbing column contributes about twice the dramatic changes in the hardness ratios of the source continuum over the two epochs. The light curves and depict that as the source got fainter and harder in the last epoch than in the previous epoch, it had higher obscuration in 2012 than in 2007.
@@
The spectrum and its light curve of the first PCA component extracted from the Suzaku HXD-PIN data are shown in Fig. 4, along with the log-eigenvalue diagram. The comparison between the time series of the XIS1 and PIN data indicates that both of them have similar changes over time, so the PCA spectrum derived from the PIN observations corresponds to the extreme hard excess of the continuum seen in of the XIS1 data. However, there is no indication of the absorbing column in the eigenvector-based multivariate analysis of the PIN data. This could mean that the absorbing material mostly blocks the energy range below 5 keV, as seen in of the XIS1 in Fig. 3. In addition, there is no PCA component containing the thermal emission lines similar to the spectrum of the XIS1, so they could be the features predominately present in the soft excess. We should also note that the PIN has an energy resolution of around 3 keV in the 10–30 keV range, which is lower than the Suzaku XIS (120 eV at 6 keV), so the PIN cannot resolve any emission line features.
Figure 4 also presents the spectra and the associated time series of the first two PCA components deduced from the NuSTAR observation. The first-order PCA spectrum shows a powerlaw-like continuum similar to that seen in of the Suzaku HXD-PIN observations. Additionally, the second PCA component may contain some emission features at energies typically associated with Ar xviii Ly 3.3 keV, Ca K 3.7 keV, and Ca K 4 keV, which are possibly present in of the XMM-Newton EPIC-pn data and of the Suzaku XIS1 observations. The light curve of the NuSTAR PCA suggests that the second component, which likely contains some line features, appears during brightening events occurring at intervals between 30 and 40 ks. However, we caution that NuSTAR’s energy resolution of eV is lower than those of the XMM-Newon EPIC-pn (80 eV) and Suzaku XIS (50 eV at 1 keV). In addition, there is no evidence for the obscuring material in the multivariate statistical analysis of the NuSTAR data, implying the absorbing material is not largely variable during the course of the NuSTAR observation similar to that seen in the Suzaku observations.
@@ 4.3 Constructing Spectra from PCA
To better evaluate the nature of the PCA components from Suzaku/XIS1, we used their corresponding photon-flux data to create the xspec-compatible spectra and the response files using the ftool program ftflx2xsp. The photon-flux spectra were reconstructed according to , where is the normalized spectrum of the th-order principal component produced by the svd function, is the mean spectrum derived from the time-sliced spectra, and is the normalized eigenvalue of the th-order principal component created via svd, representing the variability fractions. We analyzed the reconstructed spectra in the Interactive Spectral Interpretation System (isis v 1.6.2-51; Houck & Denicola, 2000) that has access to the xspec models (Arnaud, 1996).
@@
Figure 5 shows the reconstructed spectra of the first and second principal components of the Suzaku XIS1 data, which were modeled using a phenomenological model, pcfabs (diskbb apec) pcfabs (compTT zgauss), consisting of an accretion disk model (diskbb), a collisionally ionized diffuse model of the Astrophysical Plasma Emission Code (apec; Smith et al., 2001), a Comptonization model (compTT), partial covering fraction absorption components (pcfabs), and Gaussian components (zgauss). The soft excess was well reproduced using a diskbb model made of multiple blackbody components (see, e.g., Mitsuda et al., 1984; Makishima et al., 1986) and an emission spectrum produced by the collisional plasma apec model; both of them are partially covered by an absorbing column. To create phenomenologically the curvature in the hard excess, we employed a partially covered, absorbed Comptonization model (compTT) of the soft radiation in a hot plasma cloud analytically obtained by Titarchuk (1994) and Titarchuk & Lyubarskij (1995). To improve the model fit, three Gaussian components were also included, namely Fe K (6.379 keV), Fe He (6.693 keV), and Fe Ly (6.946 keV). We caution that the PCA components are built from the Suzaku data according to the temporal evolution over the two epochs, so not all the derived spectral features are physically real.
@@
Table 3 lists the best-fitting values of the parameters, obtained using the Levenberg–Marquardt minimization algorithm (Moré, 1978) and the chi-square statistic (; see Bevington & Robinson, 2003), with the uncertainties at 90% confidence derived using the isis standard function for confidence limits (conf_loop). It can be seen that the soft excess is partially covered by an absorbing column of cm-2. However, the covering fraction in the first PCA spectrum () is higher than that in the second PCA spectrum (). Moreover, the hard excess is partially absorbed by columns of and cm-2 in the first and second PCA components, respectively; both of them with a covering fraction of . This implies that there is more absorbing material in the first principal component. Interestingly, Luna et al. (2018) also found that the absorbing column in 2012 is about cm-2 higher than that in 2007 based on the XIS1 data.
@@
In addition, the curvature in the soft excess was phenomenologically reproduced using a multiple-blackbody accretion disk with a temperature of keV at the inner radius, along with a collisionally ionized soft thermal plasma emission with temperatures of 1.1 and 1.2 keV in the first and second PCA components, respectively. We should note that the inclusion of the apec collisional plasma model helps to recreate the PCA spectra, while it could not be done using only a diskbb model. In particular, Danehkar et al. (2021) previously deduced a heavily obscured, soft thermal plasma emission with a temperature of 1.3 keV using low-count Bayesian statistics. Moreover, the third PCA component of the Suzaku XIS1 data, along with the principal component of the XMM-Newton EPIC-pn data, suggests the possible presence of some emission features mostly at the soft excess, which could be associated with thermal features of a collisionally ionized plasma. So, the soft thermal apec component of – keV that appears in the spectral analysis of the PCA components might be physically related to that phenomenon.
@@
The soft component with a temperature of keV, if it exists, could be produced by the expanding winds or jets, similar to what was proposed for CH Cyg (Ezuka et al., 1998). The predicted X-ray temperature of the shock-ionized plasma created by the collision of the winds with the interstellar medium is , where is the wind velocity, is the mean molecular weight, and the hydrogen mass (e.g., Güdel & Nazé, 2009). To produce an X-ray thermal spectrum of 1.2 keV, a wind velocity of 1000 km s-1 is necessary. This is in the range of those estimated for CH Cyg (Karovska et al., 2007, 2010), which is another -type symbiotic star containing soft thermal spectra of 0.2 and 0.7 keV (Ezuka et al., 1998).
@@ 5.1 Transient Nature
The hard X-ray source IGR J12349-6434 was first detected with the IBIS instrument aboard INTEGRAL in 2003–2004 (Chernyakova et al., 2005). Its association with RT Cru was suggested by Masetti et al. (2005) and confirmed by Swift observations (Tueller et al., 2005b). Although the source was at a level of mCrab in the 20–60 keV energy band during the first detection (Chernyakova et al., 2005), it was seen 3.4 times brighter at 13 mCrab in the 18–40 keV energy band in 2012 (Sguera et al., 2012), followed by a flux level of mCrab in 2015 (Sguera et al., 2015). The ASAS and AAVSO optical light curves also indicated that RT Cru became brighter from 13.5 to 11.3 mag between 1998 and 2001; there was then a gradual decrease in the brightness to 12.1 mag in 2006, and later an increase to 11.8 mag in 2009 and 11.5 mag in 2012, followed by a decrease to 12.6 mag in 2017 (Luna et al., 2018). The Swift/BAT survey revealed that the source reached its highest X-ray brightness between 2011 and 2012 coincident with the optical peak in 2012 (Luna et al., 2018). As seen in Fig. 1, the Suzaku hardness ratios exhibited harder X-rays in 2012 than in 2007. According to the spectral fitting results by Luna et al. (2018), this transition could be related to a later increase in the column density of the absorbing material, as well as a rise in the X-ray brightness owing to higher accretion rates.